### v2

# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

Abstract
Electron-phonon coupling (EPC) is key for understanding many properties of materials such as superconductivity and electric resistivity. Although first principles density-functional-theory (DFT) based EPC calculations are used widely, their efficacy is limited by the accuracy and efficiency of the underlying exchange-correlation functionals. These limitations become exacerbated in complex $d$- and $f$-electron materials, where beyond-DFT approaches and empirical corrections, such as the Hubbard $U$, are commonly invoked. Here, using the examples of CoO and NiO, we show how the efficient r2SCAN density functional correctly captures strong EPC effects in transition-metal oxides without requiring the introduction of empirical parameters. We also demonstrate the ability of r2SCAN to accurately model phonon-mediated superconducting properties of the main group compounds (e.g., MgB2), with improved electronic bands and phonon dispersions over those of traditional density functionals. Our study provides a pathway for extending the scope of accurate first principles modeling of electron-phonon interactions to encompass complex $d$-electron materials.
IntroductionElectron-phonon coupling (EPC) is fundamental to understanding a range of phenomena in condensed matter physics and materials science Giustino (2017); Grimvall (1981), including phonon-mediated superconductivity Pellegrini and Sanna (2024), electrical resistivity in metals, carrier lifetime and mobility in semiconductors, and the temperature-dependent behavior of electronic structures Engel et al. (2020, 2022). The development of Kohn-Sham (KS) density functional theory (DFT) has enabled the quantitative prediction of EPCs through first-principles calculations, which have advanced significantly over the last few decades Giustino (2017). The two main methods that exist for calculating EPC matrix elements are density functional perturbation theory (DFPT) Baroni et al. (2001) and the finite-difference method Shang and Yang (2023); Monserrat (2018).
DFPT, often combined with maximally localized Wannier functions Giustino et al. (2007); Zhou et al. (2021a), provides an efficient means of calculating EPC matrix elements across the entire Brillouin Zone (BZ). However, it is typically limited to traditional density functional approximations (DFAs) that are not explicitly dependent on KS orbitals, such as the local-density approximation (LDA) and generalized-gradient approximations (GGAs). The finite-difference approach offers greater flexibility, allowing the inclusion of orbital-dependent density functionals and integration with advanced electronic structure solvers that consider correlation effects beyond the mean-field description Monserrat (2018); Antonius et al. (2014); Yin et al. (2013), but it often requires large supercells, making it computationally intensive.
For complex materials with $d$- and $f$-electrons, LDA and GGAs often fail to accurately capture their electronic structures Furness et al. (2018); Lane et al. (2018); Zhang et al. (2020a, 2024) and phonon properties Ning et al. (2024, 2022, 2023); Zhou et al. (2021b), resulting in inaccurate predictions for EPCs. While methods like DFPT+$U$ can improve accuracy Zhou et al. (2021b); Floris et al. (2020), they rely on parameterizations specific for $d$- and $f$-electrons. More precise approaches, such as combining GW with DFPT Li et al. (2019, 2021), further enhance EPC predictions but involve substantial computational costs. This underscores the need for efficient methods that can deliver accurate EPC predictions without the computational burden of beyond-DFT methods.
Meta-GGAs like the strongly-constrained and appropriately-normed (SCAN)  Sun et al. (2015) density functional and its variant r2SCAN Furness et al. (2020), which are orbital dependent and scale computationally similar as LDA and GGAs, have shown promise in providing accurate descriptions of total energies and electronic structures across diverse systems Sun et al. (2015); Furness et al. (2020); Sun et al. (2016). These functionals have proven effective in materials ranging from liquid water Chen et al. (2017) to metal oxides Gautam and Carter (2018), and complex materials with $d$- and $f$-electrons Zhang et al. (2020a, 2022), like the cuprates Lane et al. (2018); Furness et al. (2018); Zhang et al. (2020b) and nickelates Zhang et al. (2021); Lane et al. (2023); Zhang et al. (2024). Recent work also indicates that SCAN/r2SCAN can accurately calculate phonon dispersions of cuprates Zhang et al. (2020a); Ning et al. (2023) and NiO Ning et al. (2022). However, their performance on EPC predictions remains unexplored.
[…]
Here, we apply the r2SCAN meta-GGA within the finite-difference approach Engel et al. (2020, 2022) to calculate EPCs of transition metal compounds including cobalt oxide (CoO) and nickel oxide (NiO), as well as a main-group conventional superconductor, magnesium diboride (MgB2). We show that r2SCAN correctly captures strong EPC effects in CoO and NiO, significantly improving over the standard Perdew-Burke-Ernzerhof (PBE) GGA without requiring the introduction of empirical parameters. At the same time, r2SCAN achieves high accuracy in modeling phonon-mediated superconducting properties of MgB2.
Results and Discussion
Evaluation of EPCs in CoO and NiOCoO and NiO are two fundamental transition metal oxides that pose challenges for DFT Zhang et al. (2020a); Trimarchi et al. (2018). In our simulations, we use the low-temperature antiferromagnetic (AFM) phase of CoO and NiO, characterized by ferromagnetic planes along the [111] direction that alternate in magnetization. Upon relaxation, the cubic rock-salt structures of CoO and NiO exhibit slight distortions due to the AFM order. Experimentally, CoO and NiO are polar semiconductors, which exhibit long-range dipole-dipole interactions arising from the polarization of collective ionic movements. These interactions cause splittings between longitudinal optical (LO) and transverse optical (TO) phonon modes at the $\Gamma$ point—a phenomenon known as LO-TO splitting. To account for this effect, the nonanalytic contributions needed for calculating phonon dispersions and EPCs must be included Baroni et al. (2001); Giannozzi et al. (1991), which depend on the Born effective charges ($Z^{*}$) and the ion-clamped macroscopic dielectric tensor ($\epsilon^{\infty}$).
Figure 1(a) presents the phonon dispersions of CoO calculated using r2SCAN and PBE, along with the EPC matrix elements from r2SCAN. PBE predicts unphysical negative phonon frequencies along the $M$-$F$-$\Gamma$ and $T$-$K$-$L$ paths, indicating lattice instability in CoO, as reported in previous studies Floris et al. (2020); Zhou et al. (2021b). Consequently, PBE is unsuitable for calculating the EPC in CoO. The PBE+$U$ method is often applied to eliminate these negative frequencies, allowing DFPT+$U$ to provide accurate EPC predictions Floris et al. (2020); Zhou et al. (2021b); Wdowik and Parlinski (2007). Remarkably, r2SCAN resolves these lattice instabilities without requiring any $U$ parameters, as shown in Fig. 1(a). Furthermore, a comparison of phonon dispersions calculated by r2SCAN and PBE reveals that r2SCAN shifts phonon frequencies—particularly for the optical modes—upward, achieving better agreement with experimental results, as shown in Figs. 1(a) and S2 of the Supplementary Materials (SM).
In Fig. 1(a), our r2SCAN calculations also reveal significant LO-TO splitting in CoO. Note that PBE doesn’t capture this LO-TO splitting because it incorrectly predicts a metallic ground state for CoO [Fig. S1 (b)]. The predicted values of $\omega_{\text{LO}}=(59.4,\,62.6)$ meV and $\omega_{\text{TO}}=(38.5,\,45.7)$ meV from r2SCAN are close to experimental measurements [see Fig. S2]. Additionally, we observe notable splitting between the TO1 and TO2 modes along the $F-\Gamma-T$ direction [Fig. 1(a)]. This splitting primarily results from the structural distortion induced by the AFM order, which breaks the equivalence along the [111] directions ($\Gamma-T$), as noted in previous studies Floris et al. (2020, 2011). A similar, though smaller, LO mode splitting is observed at the $\Gamma$ point. Similar results were obtained using PBE+$U$ calculations Floris et al. (2020), although the top two optical modes along the $F-\Gamma-T$ direction are reversed in our r2SCAN results compared to PBE+$U$, likely due to the larger Born effective charge $Z^{*}$ (2.25) and ion-clamped macroscopic dielectric function $\varepsilon^{\infty}$ (6.17) values predicted by PBE+$U$ Floris et al. (2020).
CoO NiO
[…]
PBE 0 2.36 0.81 2.38 1.30 20.70 19.67
r2SCAN 0.85 1.83 2.58 5.48 5.44 2.97 2.23 1.58 7.52 7.10
Expt. 2.8 Pratt and Coelho (1959) 2.06 Sakurai et al. (1968), 1.78 Gielisse et al. (1965) 3.8 Herrmann-Ronzaud et al. (1978) 5.3 Gielisse et al. (1965) 4.0 Hüfner et al. (1984), 4.3 Sawatzky and Allen (1984) 2.2 Wang et al. (2010) 1.9 Cheetham and Hope (1983) 5.7 Gielisse et al. (1965)
Table 1: The calculated band gap ($E_{g}$), diagonals of the Born effective charge tensor ($Z^{*}$), local magnetic moments (m) and ion-clamped macroscopic dielectric constants in the RPA ($\varepsilon^{\infty}_{\text{RPA}}$) and including $f_{xc}$ ($\varepsilon^{\infty}$) of CoO and NiO for the PBE and r2SCAN functionals, as well as measured experimental values.
Figure 1(a) also shows that our r2SCAN calculations reveal substantial EPC contributions from both LO and TO modes near the $\Gamma$ point, with maximum EPC matrix elements [g defined in Eq. (1)] exceeding $10^{4}$ meV. This result aligns with previous DFPT+$U$ findings Zhou et al. (2021b) quantitatively. In transition metal oxides like CoO and NiO, the polar nature of the bonds leads to strong interactions between electrons and LO and TO phonons via the Fröhlich mechanism Fröhlich (1954). These interactions become particularly significant as the wave vector $\mathbf{q}\rightarrow 0$, causing a divergence in the EPC matrix elements for the LO and TO modes [Fig. 1 (a)]. These interactions are often substantial enough to form large polarons—quasiparticles composed of an electron or hole coupled with lattice distortions—these play a crucial role in charge transport and the dynamic behavior of electrons in polar materials. It is a significant advantage of r2SCAN to capture the Fröhlich interaction in polar materials without relying on empirical parameters Zhou et al. (2021b, 2018); Zhou and Bernardi (2016).
The phonon dispersions of NiO calculated using r2SCAN and PBE are shown in Fig. 1(b), where a similar phonon softening trend from r2SCAN to PBE is observed, as seen in CoO. Additionally, r2SCAN accurately captures the LO-TO mode splitting in NiO due to the inclusion of the nonanalytic term. Similar to CoO, r2SCAN calculations reveal a clear splitting of the TO1 and TO2 modes at the $\Gamma$ point in NiO, though the LO mode splitting is comparatively smaller, as discussed in Ref. Floris et al. (2011). Notably, the TO mode splitting in NiO is smaller than in CoO [Fig.1(b)], which can be attributed to the smaller local magnetic moments of NiO [Table 1], as pointed out in the previous study Floris et al. (2020). Importantly, the phonon characteristics computed with r2SCAN show strong consistency with those obtained using PBE+$U$, as reported in Ref. Floris et al. (2011).
Similar to CoO, strong EPC contributions from the LO and TO modes in NiO are also predicted by r2SCAN [Fig.1(b)]. In addition, to illustrate the differences between PBE and r2SCAN in EPC predictions, we plot the EPC matrix elements of NiO for the highest optical mode along the high-symmetry $\mathbf{q}$-path for each functional in Fig.1(c). The EPC matrix elements are significantly enhanced by r2SCAN compared to PBE, particularly around the $\Gamma$ point, with an approximate increase of 2.5, as shown in Fig. 1(c).
Why does r2SCAN yield superior EPC predictions in CoO and NiO? The results above clearly demonstrate that r2SCAN significantly improves the accuracy of EPC descriptions for CoO and NiO over PBE. Our analysis indicates that the superior performance of r2SCAN over PBE can largely be attributed to the more exact constraints that r2SCAN satisfies Sun et al. (2015, 2016); Furness et al. (2020). Additionally, for transition and rare-earth metal compounds, r2SCAN exhibits reduced a self-interaction error (SIE) compared to PBE, significantly improving the description of their electronic structures [Fig. 2 and S1] Zhang et al. (2020b, 2022). The significant improvement of r2SCAN over PBE in describing the phonon characteristics of NiO, particularly for the optical modes, has also been attributed to the reduction of the SIE Ning et al. (2022). SIE arises from the imperfect cancellation of the spurious classical Coulomb self-interaction by the approximate exchange-correlation self-interaction Perdew and Zunger (1981). Since the repulsive self-Coulomb term exceeds the attractive self-exchange-correlation term, the net SIE is generally positive for semi-local functionals, causing orbitals to be underbound (with orbital energies too high) and wave functions to be excessively delocalized. This leads to electron densities that are overly spread out and easily perturbed, ultimately resulting in underestimated magnetic moments and band gaps. The reduction of the SIE in r2SCAN compared to PBE therefore results in less delocalization error, more compact $d$- and $f$-orbitals, and fewer fractional occupations, which stabilize the magnetic moments Zhang et al. (2020a, 2022) and increase the band gaps for NiO and CoO as shown in Table 1.
Table 1 also presents the calculated $Z^{*}$ and $\varepsilon^{\infty}$ values for CoO and NiO from both PBE and r2SCAN, compared with experimental data. Both $Z^{*}$ and $\varepsilon^{\infty}$ are critical quantities used in the nonanalytic contributions needed for calculating phonon dispersions and EPCs. The Born effective charge $Z^{*}$ measures the dynamic charge of an atom in a crystal, capturing how polarization changes in response to the atomic displacement and accounting for the redistribution of electronic charge under an applied electric field Ghosez et al. (1998); Gonze and Lee (1997); King-Smith and Vanderbilt (1993); Resta (1993). The ion-clamped macroscopic dielectric tensor $\varepsilon^{\infty}$ describes the long-range dielectric response of a material, considering only the electronic contribution without ionic motion, reflecting how the electronic cloud in a crystal responds to an external electric field Ghosez et al. (1998); Gonze and Lee (1997); King-Smith and Vanderbilt (1993); Resta (1993). $Z^{*}$ and $\varepsilon^{\infty}$ are expected to be overestimated due to the SIE that leads to overly delocalized and easily perturbed electron densities. This can be seen by comparing the $Z^{*}$ and $\varepsilon^{\infty}$ values calculated from r2SCAN and PBE against the experimental data. For CoO, we find that the r2SCAN predictions for $Z^{*}$ and $\varepsilon^{\infty}$ are in good agreement with experimental results Sakurai et al. (1968); Gielisse et al. (1965). However, PBE incorrectly predicts a metallic ground state for CoO as shown in Fig. S1(b), to be discussed later, resulting in the vanishing of the nonanalytic terms. For NiO, PBE significantly overestimates $\varepsilon^{\infty}$ by a factor of 3.77, whereas r2SCAN provides significant improvement, reducing the overestimation to 31.93%. The calculated $Z^{*}$ value using r2SCAN is 2.23, which is in excellent agreement with the experimental value of 2.2, while PBE overestimates $Z^{*}$ by 8.18%. These results corroborate that r2SCAN reduces SIE in comparison with PBE for NiO and CoO.
The above analysis strongly indicates that the improvement of r2SCAN over PBE on EPC matrix elements originates from the SIE reduction. In the context of the Kohn-Sham (KS) DFT Giustino (2017), the EPC matrix element is defined as:
$g_{mn,\nu}({\boldsymbol{k}},\boldsymbol{q})=\langle\psi_{m{\boldsymbol{k}+\boldsymbol{q}}}|\partial_{\boldsymbol{q}\nu}v^{\text{KS}}|\psi_{n{\boldsymbol{k}}}\rangle,$ (1)
where $\psi_{n{\boldsymbol{k}}}$ is the electronic wavefunction with the momentum $\boldsymbol{k}$ in the band $n$, and $\partial_{\boldsymbol{q}\nu}v^{\text{KS}}$ the derivative of the self-consistent KS potential with respect to a phonon of wavevector $\boldsymbol{q}$, band index $\nu$, and frequency $\omega_{\boldsymbol{q}\nu}$. We note that $\partial_{\boldsymbol{q}\nu}v^{\text{KS}}$ is calculated by the finite difference approach here. By definition $\partial_{\boldsymbol{q}\nu}v^{\text{KS}}$ is given by
$\partial_{\boldsymbol{q}\nu}v^{\text{KS}}=\partial_{\boldsymbol{q}\nu}v^{\text{en}}\ +\partial_{\boldsymbol{q}\nu}v^{\text{H}}\ +\partial_{\boldsymbol{q}\nu}v^{\text{xc}},$ (2)
where $v^{\text{en}}$ is the potential arising from electron-nuclei interactions, $v^{\text{H}}$ is the Hartree potential, and $v^{\text{xc}}$ is the exchange-correlation potential. This can be recast into Onida et al. (2002):
$\partial_{\mathbf{q}\nu}v^{\text{KS}}=(\varepsilon^{\text{Hxc}})^{-1}\partial_{\mathbf{q}\nu}v^{\text{en}}.$ (3)
We have used the definition of the ‘test electron’ dielectric matrix,
$\varepsilon^{\text{Hxc}}=1-(v^{\text{C}}+f^{\text{xc}})\chi^{0}.$ (4)
Here, $v^{\text{C}}$ is the Coulomb kernel, $f^{\text{xc}}(\mathbf{r},\mathbf{r}^{\prime})\equiv\frac{\delta^{2}E_{\text{xc}}[n]}{\delta n(\mathbf{r})\delta n(\mathbf{r}^{\prime})}$ the exchange-correlation kernel, and $\chi^{0}(\mathbf{r},\mathbf{r}^{\prime})\equiv\frac{\delta n(\mathbf{r})}{\delta v^{\text{KS}}(\mathbf{r}^{\prime})}$ the KS non-interacting linear response function.
Based on Eqs. (3) and (4), one can see that the change in the KS potential due to the motion of the nuclei originates from two contributions: (i) the change in the bare electron-nuclei potential, and (ii) the response of the electronic degrees of freedom to screen these interactions at the DFA level. Thus, improvements of EPC matrix elements within the KS DFT framework should come from two key contributions: $\chi_{0}$ and $f^{\text{xc}}$.
Figure 2: The calculated electronic structure of $G$-type AFM phase of CoO and NiO using the r2SCAN functional. (a) Orbitally-projected band structure of CoO along high-symmetry lines in the BZ, where the Co $t_{2g}$ and $e_{g}$, and O $p$ orbitals are marked with blue, red, and green colors, respectively. The side panel shows the partial density of states of the respective orbitals. (b) Same as (a), but for NiO.
The ion-clamped macroscopic dielectric constants $\varepsilon^{\infty}$ are quantities related to the head ($\mathbf{q}\rightarrow 0$) of $\varepsilon^{\text{Hxc}}$ Gajdoš et al. (2006). For a given density functional, $\varepsilon^{\infty}$ can be calculated using the Random Phase Approximation (RPA) where $f^{\text{xc}}$ is ignored, with the value denoted as $\varepsilon^{\infty}_{\text{RPA}}$ Gajdoš et al. (2006); Paier et al. (2008). Table 1 shows that the difference between $\varepsilon^{\infty}_{\text{RPA}}$ and $\varepsilon^{\infty}$ of a density functional is much smaller than the difference between $\varepsilon^{\infty}$ of different density functionals, suggesting that $f^{\text{xc}}$ has a less important effect on $\varepsilon^{\infty}$ than the $\chi_{0}$ effect of the chosen density functional. We note that r2SCAN satisfies more exact constraints than PBE does with more accurate total energies in general, suggesting that r2SCAN can have a better $f^{\text{xc}}$ Nazarov and Vignale (2011).
The KS linear response function $\chi_{0}$ describes how the electron density $n$ reacts to small changes in the Kohn-Sham potential $v^{\text{KS}}$. This linear response function is often overestimated in magnitude by SIE, which leads to an overly delocalized and easily perturbable electron density, as well as an underestimated band gap that is inversely proportional to the linear response function.
For CoO, PBE fails to open a band gap, as illustrated in Fig. S1(b). This leads to too large $\chi_{0}$, resulting in dynamical instabilities with the appearance of imaginary frequencies in the phonon calculations as shown in Figure 1(a). In contrast, due to the SIE reduction, r2SCAN predicts an indirect band gap of 0.85 eV, attributed to the splitting between the occupied $t_{2g}$ orbitals and the unoccupied $t_{2g}$ orbitals in the spin-down channel, as shown in Fig. 2(a). This undoubtedly improves $\chi_{0}$ and thus provides an improved phonon dispersion as well as strong EPC effects at the LO and TO phonon modes around the $\Gamma$ point as shown in Figure 1(a). Since a smaller gap generally correlates with a larger $\varepsilon^{\infty}$ Penn (1962); Onishi and Fu (2024), the improved gaps from r2SCAN explain our values of $\varepsilon^{\infty}$ being closer to experimental results as shown in Table. 1.
For NiO, r2SCAN predicts an indirect band gap of 2.97 eV, with the valence band maximum (VBM) at the $T$ point and the conduction band minimum (CBM) along the $T$-$K$ direction, significantly larger than the 0.8 eV band gap predicted by PBE [Fig. S1(d)] but still noticeably smaller than the 4.0$\sim$4.3 eV of experimental results [see Table 1]. PBE significantly underestimates the band gap, yielding a larger $\chi_{0}$, which, in turn, produces smaller EPC matrix elements compared to r2SCAN, as shown in Fig. 1(c).
From the above analysis, accurate EPC predictions can be expected within the KS-DFT framework if the SIE in the exchange-correlation approximation is minimized. This can be achieved by advancing up Jacob’s ladder of DFT Perdew and Schmidt (2001), moving from LDA to GGA, to meta-GGA, and then to hybrid density functionals, PZ-SIC Perdew and Zunger (1981), etc., or by applying empirical corrections to the chosen functional, such as DFT+$U$ Anisimov et al. (1991, 1993). The HSE hybrid density functional Heyd et al. (2003) has become a standard for calculating band gaps in solids due to its reduced SIE; however, its computational cost—one or two orders of magnitude higher than LDA, GGAs, and metaGGAs—limits its practicality for EPC calculations. DFT+$U$ offers an alternative to metaGGA functionals like r2SCAN, but the empirical nature of the $U$ parameters reduces its predictive power Anisimov et al. (1991, 1993).
Performance of r2SCAN in conventional superconductor MgB2Having examined the improvements that r2SCAN offers for predicting EPCs in complex compounds like CoO and NiO with $d$-electrons, we now demonstrate that r2SCAN also provides reliable predictions of the EPC strength $\lambda$ [defined in Eq. (6)] in main group compounds. Here, we revisit the prototypical conventional superconductor MgB2, a well-known phonon-mediated multiband superconductor with a critical temperature ($T_{\text{c}}$) of approximately 39 K at ambient pressure Nagamatsu et al. (2001); Chen et al. (2001); Tsuda et al. (2001). MgB2 has been extensively studied, both experimentally Nagamatsu et al. (2001); Chen et al. (2001); Tsuda et al. (2001); Baron et al. (2004); d’Astuto et al. (2007a) and theoretically Choi et al. (2002a); Margine and Giustino (2013); Kortus et al. (2001); Yildirim et al. (2001); Choi et al. (2002b).
MgB2 crystallizes in the tetragonal Bravais lattice with space group $P6/mmm$ (#191, 6/mmm), where boron atoms form honeycomb lattices separated by hexagonal layers of magnesium atoms [Fig. S3]. The optimized lattice parameters of MgB2 using r2SCAN are $a=3.063$ and $c=3.521$, which are in excellent agreement with the corresponding experimental values $a=3.086$ and $c=3.524$ Nagamatsu et al. (2001). The electronic states at the Fermi level calculated from r2SCAN are primarily derived from boron $p$-states. Specifically, two $\sigma$-bonding bands, contributed by B-$p_{x/y}$ orbitals, cross the Fermi level [Fig. 3(a)], forming two small cylindrical hole Fermi surfaces along the $\Gamma-A$ direction [Fig. 3(e)]. These $\sigma$ bands are confined to the boron planes, exhibiting a two-dimensional character. In contrast, the dispersive $\pi$-bonding bands, derived from the B-$p_{z}$ orbitals, exhibit a three-dimensional character [Fig. 3(e)], resulting in one hole and one electron Fermi surfaces near the $k_{z}=0$ and the $k_{z}=\pm\pi/c$ planes, respectively. Our findings are generally consistent with previous calculations using standard DFT Kortus et al. (2001); Yildirim et al. (2001). However, subtle differences between our r2SCAN results and those obtained using PBE are noticeable, as illustrated in Fig. 3 (a). For instance, the bandwidth predicted by r2SCAN is generally larger than that predicted by PBE. At the $M$ point, the band splitting between the upper $\pi$ band and the lower $\sigma$ band is approximately 2.6 eV with r2SCAN, compared to 2.4 eV with PBE. Notably, the r2SCAN prediction matches the experimental value of 2.6 eV Guritanu et al. (2006), which also agrees with high-level calculations such as modified HSE of 2.6 eV Yin et al. (2013) and scQPGW of 2.6 eV Yin et al. (2013).
Figure 3(c) presents the calculated phonon dispersion of MgB2 using the r2SCAN and PBE functionals. Consistent with prior studies d’Astuto et al. (2007a); Yildirim et al. (2001), r2SCAN identifies four optical phonon modes at the $\Gamma$ point: $E_{1u}$ , $A_{2u}$, $E_{2g}$, and $B_{1g}$, with frequencies of 42.6, 50.2, 74.1, and 87.9 meV, respectively, in close agreement with experimental values d’Astuto et al. (2007a). Notably, the $E_{2g}$ mode, involving in-plane stretching of B atoms, is critical for EPCs in MgB2 Choi et al. (2002b); Margine and Giustino (2013); Choi et al. (2002a), as we can see in Figs. 3 (c) and (f). This comparison between r2SCAN and PBE results [Fig. 3(c)] shows excellent agreement for acoustic modes, while r2SCAN provides a higher and more accurate prediction for the $E_{2g}$ mode frequency of 74.1 meV, matching the experimental value of 75 meV obtained from Raman measurements d’Astuto et al. (2007a); Hlinka et al. (2001); Goncharov et al. (2001). PBE predicts a lower value of 63 meV for the $E_{2g}$ mode. This discrepancy with the experimental value is often attributed to the overestimated bond lengths by PBE, resulting in softer optical phonon modes. While previous studies have linked the difference between PBE predictions and experimental values to anharmonic effects Yildirim et al. (2001); Choi et al. (2002b), recent findings suggest that MgB2 exhibits only weak anharmonicity d’Astuto et al. (2007b). This suggests that the improved results by r2SCAN arise from a better description of exchange-correlation effects.
Figure 3: Calculated electronic structure, phonon dispersion and EPC of MgB2. (a) The calculated band structure of MgB2 along high-symmetry lines in the BZ from r2SCAN (solid black lines) and PBE (dash gray lines). Orbital characters from r2SCAN are also shown, where Mg $s$, B $p_{x/y}$, and B $p_{z}$ orbitals are indicated in green, red, and blue, respectively. The splitting of $\pi$ and $\sigma$ bands from r2SCAN at the $M$ point is highlighted. Orbital projections from PBE are omitted for clarity. (b) The total and partial density of states for each orbital using r2SCAN. (c) Phonon dispersion along high-symmetry lines in the BZ, comparing r2SCAN (solid black lines) and PBE (dashed gray lines). Marker color intensity corresponds to the magnitude of the EPC strength, $\lambda_{\textbf{q}\nu}$ from r2SCAN calculations. EPC strength $\lambda_{\textbf{q}\nu}$ values from PBE calculations are not shown. Experimental phonon data from inelastic X-ray scattering (magenta solid points), extracted from Ref. d’Astuto et al. (2007a), are also included for comparison. The measured frequency of the $E_{2g}$ mode at the $\Gamma$ point from Raman spectroscopy is 75 meV d’Astuto et al. (2007a); Hlinka et al. (2001); Goncharov et al. (2001), represented by a blue cross. (d) Calculated Eliashberg spectral functions $\alpha^{2}F(\omega)$ and cumulative EPC strength $\lambda$ for MgB2, using r2SCAN (red lines) and PBE (green lines). A k-mesh of $100\times 100\times 100$ and a q-mesh of $50\times 50\times 50$ were used for the BZ sampling. (e) Left panel: The calculated three-dimensional Fermi surface of MgB2 obtained using the Wannier model based on r2SCAN calculations with a $200\times 200\times 200$ k-mesh. Middle and right panels: The Fermi surface projections on the $k_{z}=0$ and $k_{z}=\pi/c$ planes, respectively. (f) Distribution of $k$-resolved EPC strength $\lambda_{\boldsymbol{k}}=\sum_{mn,\nu\boldsymbol{q}}|g_{mn,\nu}({\boldsymbol{k}},\boldsymbol{q})|^{2}\delta(\varepsilon_{m\textbf{k}+\textbf{q}}-\varepsilon_{\text{F}})\delta(\varepsilon_{n\textbf{k}}-\varepsilon_{\text{F}})\times(2/\omega_{\textbf{q}\nu})$ from r2SCAN across the $k_{x}-k_{y}$ plane at $k_{z}=0$ (left panel) and $k_{z}=\pi/c$ (right panel).
Figure 3(d) shows the calculated isotropic Eliashberg function $\alpha^{2}F(\omega)$ of MgB2, based on Eq. (9), using the r2SCAN (red lines) and PBE (blue lines) functionals. Our PBE results reproduce the dominant peak around 60 meV, primarily contributed by the E2g mode, consistent with previous findings Margine and Giustino (2013). A second peak appears around 91 meV, slightly higher than previous results obtained with the LDA functional Margine and Giustino (2013). The calculated $\alpha^{2}F(\omega)$ using the PBE functional yields an EPC strength $\lambda$ of approximately 0.70, close to the earlier value of 0.74 reported for the LDA functional Margine and Giustino (2013). The r2SCAN results, shown as a solid red line, exhibit similarities to the PBE results in the range of 0–50 meV, since the predicted phonon dispersions from both functionals are nearly identical for low-frequency phonons [Fig. 3 (c)]. However, the dominant peaks from the r2SCAN calculations appear in the range of 63–75 meV, which are much broader than those from the PBE results, reflecting the larger phonon bandwidth of the $E_{2g}$ mode predicted by r2SCAN. The second peak at 91 meV is also captured in our r2SCAN results. These results also reflect the softening of the optical phonon modes in PBE. The broader distribution of the dominant peak in $\alpha^{2}F(\omega)$ observed with r2SCAN results in an EPC strength $\lambda$ of 0.73, slightly larger than the PBE value of 0.70. According to Eqs. (10) and (11), the estimated $\omega_{log}$ is $\sim$58.1 meV, and the predicted $T_{\text{c}}$ ranges from 22.4$\sim$30.3 K, with the adjustable parameter $\mu^{*}$ set between 0.12 and 0.08, consistent with previous calculations Margine and Giustino (2013).
Based on the above analysis, our results demonstrate that the r2SCAN functional provides a reliable description of electronic structure, phonon dispersion, and EPC in MgB2 compared to PBE. While previous studies have shown that MgB2 is a multi-gap superconductor Tsuda et al. (2001); Chen et al. (2001); Choi et al. (2002a, b); Giustino et al. (2007); Margine and Giustino (2013); Floris et al. (2005), solving the superconducting gap equation is beyond the scope of the present work.
ConclusionCalculating EPCs for complex materials with $d$- and $f$-electrons has long been a challenge for first-principles methods due to the need for both accuracy and efficiency. In this study, we demonstrate that the efficient r2SCAN density functional significantly improves EPC predictions compared to the standard PBE density functional for transition-metal compounds such as CoO and NiO, alongside better descriptions of their electronic structures and phonon dispersions. Due to SIE, PBE incorrectly predicts CoO as metallic with imaginary phonon frequencies, rendering it unsuitable for EPC calculations. In contrast, r2SCAN reduces SIE, opens the band gap, stabilizes the phonon dispersion, and predicts strong EPCs for optical modes. Analysis of the dielectric screening tensors for CoO and NiO also indicates that these improvements stem from the reduction of SIE in r2SCAN compared to PBE.
Additionally, for the main-group compound MgB2, r2SCAN provides improved predictions of the band structure and phonon dispersion compared to PBE. Notably, r2SCAN predicts a phonon frequency for the critical E2g mode that closely matches Raman measurements, whereas PBE significantly underestimates this value. The resulting EPC strength $\lambda$ and isotropic Eliashberg spectral function $\alpha^{2}F(\omega)$ calculated using r2SCAN show good agreement with prior studies.
Our findings highlight the general efficacy of the r2SCAN functional in accurately predicting EPCs across a wide range of materials. The study underscores the potential of advanced functionals like r2SCAN to enhance the predictive power of first-principles EPC calculations, enabling deeper and more accurate insights into the physical properties of materials with strong electron-electron and electron-phonon interactions.
MethodsFirst-principles calculations were performed using the projector-augmented-wave (PAW) method Kresse and Joubert (1999), as implemented in the Vienna ab initio simulation package (VASP) Kresse and Hafner (1993); Kresse and Furthmüller (1996). Both the r2SCAN and PBE density functionals were employed in our calculations. For calculating the EPC, we follow the Wannier-interpolation procedure outlined in Ref. Engel et al. (2020) while using the all-electron EPC matrix elements defined in Ref. Engel et al. (2022); Chaput et al. (2019). The Wannier representation of the electronic band structure was obtained using Wannier90 Mostofi et al. (2014) with the help of the VASP2WANNIER90 interface. We calculate the EPC strength $\lambda$ and the isotropic Eliashberg spectral function $\alpha^{2}F$ based on the following equations.
Within the Migdal approximation Migdal (1958), the imaginary part of the phonon self-energy is expressed as
$\Pi^{\prime\prime}_{\boldsymbol{q}\nu}=\text{Im}\sum_{m,n,\boldsymbol{k}}w_{\boldsymbol{k}}|g_{mn,\nu}({\boldsymbol{k},\boldsymbol{q}})|^{2}\frac{f_{n\boldsymbol{k}}-f_{m{\boldsymbol{k}+\boldsymbol{q}}}}{\epsilon_{m{\boldsymbol{k}+\boldsymbol{q}}}-\epsilon_{n\boldsymbol{k}}-\omega_{\boldsymbol{q}\nu}-i\eta},$ (5)
where $f_{n{\boldsymbol{k}}}$ is the Fermi occupation of the single-particle state $|\psi_{n\boldsymbol{k}}\rangle$ associated with the eigenvalue $\epsilon_{n\boldsymbol{k}}$, $w_{\boldsymbol{k}}$ are the weights of the dense ${\boldsymbol{k}}$-points, and $\eta$ indicates a real positive infinitesimal. We note here that the spin index is omitted for simplicity. The EPC strength $\lambda$ is the full BZ average of the mode-resolved $\lambda_{\boldsymbol{q}\nu}$
$\lambda=\sum_{\boldsymbol{q}\nu}w_{\boldsymbol{q}}\lambda_{\boldsymbol{q}\nu},$ (6)
with $w_{\boldsymbol{q}}$ being the weights of the dense $\boldsymbol{q}$-points. In our implementations, the double delta approximation Marini et al. (2024) is adopted to calculate the imaginary part of the phonon self-energy, under which the Eq. (5) can be rewritten as follows:
$\Pi^{\prime\prime}_{\boldsymbol{q}\nu}=\text{Im}\sum_{m,n,\boldsymbol{k}}w_{\boldsymbol{k}}|g_{mn,\nu}({\boldsymbol{k},\boldsymbol{q}})|^{2}\delta(\epsilon_{n{\bf k}})\delta(\epsilon_{m{\bf k}+{\bf q}})$ (7)
Under the double delta approximation, the $\lambda_{\boldsymbol{q}\nu}$ can be expressed as:
$\lambda_{\boldsymbol{q}\nu}=\frac{1}{\pi N_{\text{F}}}\frac{\Pi^{\prime\prime}_{\boldsymbol{q}\nu}}{\omega^{2}_{\boldsymbol{q}\nu}},$ (8)
where $N_{\text{F}}$ is the density of states at the Fermi level. Using this mode-resolved $\lambda_{\boldsymbol{q}\nu}$ and the phonon frequencies $\omega_{\boldsymbol{q}\nu}$, the isotropic Eliashberg spectral function, $\alpha^{2}F$, can be calculated by
$\alpha^{2}F(\omega)=\frac{1}{2}\sum_{\boldsymbol{q}\nu}w_{\boldsymbol{q}}\omega_{\boldsymbol{q}\nu}\lambda_{\boldsymbol{q}\nu}\,\delta(\omega-\omega_{\boldsymbol{q}\nu}).$ (9)
The phonon-mediated superconducting transition temperature, $T_{\text{c}}$, is evaluated with the McMillian-Allen-Dynes formula McMillan (1968),
$T_{\text{c}}=\frac{\omega_{\text{log}}}{1.2}\exp\left[{\frac{-1.04(1+\lambda)}{\lambda(1-0.62\mu^{*})-\mu^{*}}}\right],$ (10)
where the logarithmic average frequency $\omega_{\text{log}}$ can be calculated from the following equation Allen and Dynes (1975); Grimvall (1981)
$\omega_{\text{log}}=\exp\left[\frac{2}{\lambda}\int^{\infty}_{0}\text{d}\omega\frac{\log\,\omega}{\omega}\alpha^{2}F(\omega)\right].$ (11)
CoO and NiO All the calculations were performed within the density functional theory framework with projector-augmented wave potentials Kresse and Joubert (1999) using the VASP Kresse and Hafner (1993); Kresse and Furthmüller (1996). An initial $G$-type antiferromagnetic (AFM) structure with 4 atoms was used for structural relaxation. A high-energy cutoff of 600 eV was applied to truncate the plane-wave basis set. Crystal structures and ionic positions were fully optimized with a force convergence criterion of 0.001 eV/Å for each atom and a total energy tolerance of $10^{-8}$ eV. A $\Gamma$-centered k-point mesh of $15\times 15\times 15$ was employed for the structural relaxation. For phonon calculations, a $4\times 4\times 4$ supercell with a $\Gamma$-centered $k$-point mesh of $3\times 3\times 3$ was used, employing the finite displacement method. A Gaussian smearing with a small $\sigma$ of 0.01 was applied in our phonon calculations. Transition metal (Ni and Co) $d$ states and oxygen $p$ states were included in constructing the Wannier functions.
MgB2The calculations were performed within the density functional theory framework with projector-augmented wave potentials Kresse and Joubert (1999) using the VASP Kresse and Hafner (1993); Kresse and Furthmüller (1996). A high-energy cutoff of 520 eV was applied to truncate the plane-wave basis set. Crystal structures and ionic positions were fully optimized using a force convergence criterion of 0.001 eV/Å for each atom, along with a total energy tolerance of $10^{-8}$ eV. To relax the structures, we adopted a $\Gamma$-centered k-point mesh of $13\times 13\times 12$ to sample the primitive bulk BZ. For phonon calculations, a $6\times 6\times 6$ supercell with a $\Gamma$-centered k-point mesh of $7\times 7\times 9$ was employed using the finite displacement method. A Gaussian smearing with a small $\sigma$ of 0.01 was applied in our phonon calculations. Mg $p$ and $s$ states, along with B $p$ and $s$ states, were included in constructing the Wannier functions.